# The Unbounded Denominators Conjecture

### v3

@@ 5.2. Geometry of $\Gamma_{N}$ and a uniform growth estimate of $F_{N}$
The idea of the proof is to use the symmetry of $\mathbf{C}\smallsetminus\mu_{N}$ and the action of the Fuchsian group $\widetilde{\Gamma}_{N}$ to reduce the question to the study of the asymptotic of $\widetilde{F}_{N}$ near the cusp $\tau=i\infty$ and we study this asymptotic using the explicit description of the inverse of $\widetilde{F}_{N}$ given in Lemma 5.1.14.
@@
The group $\widetilde{\Gamma}_{N}$ is the free group on $N$ generators generated by $\widetilde{t}_{N}$ and its conjugates by powers of $\widetilde{r}_{N}$ in (5.1.3). A fundamental domain $\Omega\subset D(0,1)$ for $\Gamma_{N}$ is given by the region with $2N$ cusps given by half-integer powers of $\zeta_{N}=\exp(2\pi i/N)$ and bounded by geodesics connecting adjacent cusps.
@@
Since $\widetilde{F}_{N}$ is the universal covering map to $\mathbf{C}\smallsetminus\mu_{N}$, then the Fuchsian group $\widetilde{\Gamma}_{N}$ is generated by the stabilizers of the cusps $c$ with $\widetilde{F}_{N}(c)\in\mu_{N}$. If we denote the generator of the stabilizer of $i\infty$ by
@@
$\{z\in D(0,1)\,:\,d(gz,0)\geq d(z,0)\quad\forall g,g^{-1}\in\{r_{N}^{k}\,t\,r_{N}^{-k}\},k=0,1,\ldots,N-1\}.$
Here, $d$ is the hyperbolic distance in $D(0,1)$. The region in $D(0,1)$ such that $d(gz,0)\geq d(z,0)$ and $d(g^{-1}z,0)\geq d(z,0)$ for $g=r_{N}^{k}\cdot t\cdot r_{N}^{-k}$ is the region bounded by two geodesics starting at $\zeta_{N}^{k}$ going in opposite directions and intersecting the boundary at $\zeta_{N}^{k}e^{\pm i\theta}$ where $c_{N}=2\cot(\theta/2)$. There are exactly $2N$ such arcs corresponding to the $N$ generators and their inverses. In particular, if $\theta<\pi/N$ is too small, the fundamental region will have infinite volume, whereas if $\theta>\pi/N$ is too big, then the Dirichlet domain will only contain at most $N$ cusps. Since $\mathbf{H}/\widetilde{\Gamma}_{N}$ has $N+1$ cusps and $\widetilde{\Gamma}_{N}$ has finite covolume, we must have $c_{N}=2\cot(\pi/2N)$ and these geodesics intersecting at $\zeta_{N}^{k+1/2}$, $k=0,\ldots,N-1$. ∎
@@
where $\psi$ is the digamma function as in Lemma 5.1.18, and the implicit coefficient depends on $M_{0}$ but not on $N,M$.
@@
where the implied constants are uniform in $N$, from which the result follows (using that $|\log{x}|\gg N$). ∎
@@
once $N\gg_{\epsilon,M_{0}}1$, where the implicit coefficient depends only on $M_{0}$ and $\epsilon$.
@@
If we choose $\varepsilon_{0}:=\epsilon/2$, then for $N\gg_{\epsilon,M_{0}}1$ this lower bound clearly exceeds the requisite
$\frac{2N^{2}M(1-\epsilon)}{\pi^{2}}.$
@@
Set $S(M,N):=\{z\in D(0,1)\,:\,|F_{N}^{N}(z)-1|<e^{-MN}\}$ and $M:=\frac{N}{1-r}$. Since $M\geq 2$, we may take $M_{0}=2$ and $\epsilon=1/2$ in Lemma 5.2.11 and conclude that for $N\gg 1$ (with absolute constant here), we have
@@
By Lemma 5.2.5, the set $\{z\in D(0,1)\,:\,|F_{N}(z)|>e^{M}+1\}$ is contained in $sS(M,N)$, which is contained in $D(0,1)\smallsetminus\overline{D(0,r)}$ by the above argument for $N\gg 1$. Thus we conclude that
@@
A more refined bound is proved in Kraus–Roth [KR16, Theorems 1.2 and 1.10]. On the other hand, one can push our method further and prove, with rather more work but uniformly in $N\in\mathbf{N}$ and $M\in[1,\infty)$, that the supremum region $|F_{N}|<e^{M}$ is simply connected of conformal radius $1-O(M^{-2}N^{-3})$ from the origin; this is a sharp estimate. But taking for $\varphi$ in Theorem 2.0.5 the pullback of $F_{N}$ by the Riemann map of some such region $|F_{N}|<e^{M}$, and ignoring thus the fine savings from the integrated bound (2.0.7) as opposed to the supremum, would only lead to an $O(N^{4})$ holonomy rank bound in place of our requisite logarithmically inflated bound $O(N^{3}\log{N})$. In the next section we will see how to make the full use of the integrated holonomy bound, and use Nevanlinna’s value distribution theory to supply our final piece of the proof of the unbounded denominators conjecture.
@@ 6. Nevanlinna theory and uniform mean growth near the boundary
for any universal covering map $F:D(0,1)\to\mathbf{C}\smallsetminus\{a_{1},\ldots,a_{N}\}$ based at $F(0)=0$ (see also the discussion in Nevanlinna [Nev70, page 272]), however this is only asymptotically in $r\to 1^{-}$ for given punctures $\{a_{i}\}$ whereas we need a uniformity in both $r$ and $N$. It is at the point (6.2.3) of the explicit partial fraction coefficients that our argument below makes a critical use of the special feature of the target set $\mu_{N}\cup\{\infty\}$ of omitted values.
@@
we have uniformly in $N\in\mathbf{N}$ and $r\in(0,1)$ the mean growth bound
@@
with some (effectively computable) absolute constant implicit coefficient.
@@ 6.1.1. The Nevanlinna characteristic
where, in general for a meromorphic mapping $f:D(0,1)\to\mathbf{P}^{1}$, we let $\mathrm{ord}_{\rho}^{-}(f):=\mathrm{ord}^{+}(1/f)=\max(0,\mathrm{ord}(1/f))$ is the pole order (if $\rho$ is a pole, and 0 if $f$ is holomorphic at $\rho$).
@@
For every meromorphic function $f:D(0,1)\to\mathbf{P}^{1}$ regular at 0 (that is: $f(0)\neq\infty$), and every $r\in(0,1)$, we have
@@ 6.1.2. The lemma on the logarithmic derivative
The lemma on the logarithmic derivative—a strong explicit form of which is cited in (6.1.12) below—is the centerpiece of R. Nevanlinna’s original analytic proof of his second main theorem of value distribution theory. The logarithmic error feature of this sharp upper bound on the proximity function of a logarithmic derivative enables us to derive Theorem 6.0.1 from the relatively crude supremum growth bound in Lemma 5.2.17.
@@
Differentiation in the integrand of (6.1.8) gives the reproducing formula
(6.1.9) $\frac{g^{\prime}(z)}{g(z)}=\int_{|w|=R}\frac{2w}{(w-z)^{2}}\log{|g(w)|}\,\mu_{\mathrm{Haar}}(w),\quad\forall z\in D(0,R).$
@@
$\int_{|z|=r}\Big|\frac{g^{\prime}(z)}{g(z)}\Big|\,\mu_{\mathrm{Haar}}\leq 2R\int_{|z|=r}\int_{|w|=R}|w-z|^{-2}\,\big|\log{|g(w)|}\big|\,\mu_{\mathrm{Haar}}(w)\,\mu_{\mathrm{Haar}}(z)$
$=2R\int_{|w|=R}\Big(\int_{|z|=r}|w-z|^{-2}\,\mu_{\mathrm{Haar}}(z)\Big)\,\big|\log{|g(w)|}\big|\,\mu_{\mathrm{Haar}}(w)$
$=\frac{2R}{R^{2}-r^{2}}\int_{|w|=R}\big|\log{|g(w)|}\big|\,\mu_{\mathrm{Haar}}(w)$
$=\frac{2R}{R^{2}-r^{2}}\Big(m(R,g)+m(R,1/g)\Big)=\frac{4R\,m(R,g)}{R^{2}-r^{2}},$
@@
$\int_{|z|=r}\log^{+}{\Big|\frac{g^{\prime}}{g}\Big|}\,\mu_{\mathrm{Haar}}\leq\mu_{\mathrm{Haar}}(E)\log^{+}\Big(\frac{1}{\mu_{\mathrm{Haar}}(E)}\int_{E}\Big|\frac{g^{\prime}(z)}{g(z)}\Big|\,\mu_{\mathrm{Haar}}(z)\Big)$
$\leq\log^{+}\int_{|z|=r}\Big|\frac{g^{\prime}(z)}{g(z)}\Big|\,\mu_{\mathrm{Haar}}(z)+\sup_{t\in(0,1]}\big\{t\log{(1/t)}\big\}$
$\leq\log^{+}{\Big\{\frac{4R\,m(R,g)}{R^{2}-r^{2}}\Big\}}+\frac{1}{e}\leq\log^{+}{\Big\{\frac{m(R,g)}{r}\frac{R}{R-r}\Big\}}+\log{2}+\frac{1}{e},$
@@
and proved that it is essentially best-possible in form apart for the value of the free numerical constant $5.8501$ (that has since been somewhat further reduced in the literature, see Benbourenane–Korhonen [BK01]). The paper of Hinkkanen [Hin92] and the books of Cherry–Ye [CY01] and Ru [Ru21] discuss the implications to the structure of the error term in Nevanlinna’s second main theorem, mirroring Osgood and Vojta’s dictionary to Diophantine approximation and comparing to Lang’s conjecture modeled on Khinchin’s theorem.
@@ 6.2.1. Equivalence of bounds for different $p(x)$
$T(r,p\circ F_{N})=m\Big(r,1+\frac{1}{F_{N}^{N}-1}\Big)\geq m\Big(r,\frac{1}{F_{N}^{N}-1}\Big)-\log{2}$
$=T\Big(r,\frac{1}{F_{N}^{N}-1}\Big)-\log{2}=T(r,F_{N}^{N}-1)-\log{2}$
(6.2.2) $\geq T(r,F_{N}^{N})-2\log{2}=N\,T(r,F_{N})-\log{4},$
@@ 6.2.3. Two corollaries of the lemma on the logarithmic derivative
$m\Big(r,\frac{f^{\prime}}{f}\Big)\ll\log^{+}{T\Big(\frac{1+r}{2},f\Big)}+\log{\frac{e}{1-r}}$
$=\log^{+}{m\Big(\frac{1+r}{2},1-F_{N}^{N}\Big)}+\log{\frac{e}{1-r}}$
$\ll\log^{+}{m\Big(\frac{1+r}{2},F_{N}^{N}\Big)}+\log{\frac{e}{1-r}}$
$\ll\log^{+}{m\Big(\frac{1+r}{2},F_{N}\Big)}+\log{\frac{N}{1-r}}$
$\ll\sup_{|z|=(1+r)/2}\log^{+}{\log{|F_{N}|}}+\log{\frac{N}{1-r}},$
@@
$m\Big(r,\frac{F_{N}}{F_{N}^{\prime}}\Big)=m\Big(r,\frac{F_{N}^{\prime}}{F_{N}}\Big)+N\Big(r,\frac{F_{N}^{\prime}}{F_{N}}\Big)-N\Big(r,\frac{F_{N}}{F_{N}^{\prime}}\Big)-\log{c(F_{N}^{\prime}/F_{N},0)}$
$=m\Big(r,\frac{F_{N}^{\prime}}{F_{N}}\Big)+N\Big(r,1/F_{N}\Big)-N\Big(r,F_{N}\Big)-N\Big(r,1/F_{N}^{\prime}\Big)+N\Big(r,F_{N}^{\prime}\Big)$
(6.2.10) $=m\Big(r,\frac{F_{N}^{\prime}}{F_{N}}\Big)+N\Big(r,1/F_{N}\Big)-N\Big(r,1/F_{N}^{\prime}\Big)=m\Big(r,\frac{F_{N}^{\prime}}{F_{N}}\Big)+N\Big(r,1/F_{N}\Big).$
@@
$m\Big(r,\frac{F_{N}}{F_{N}^{\prime}}\Big)=m\Big(r,\frac{F_{N}^{\prime}}{F_{N}}\Big)+N\Big(r,1/F_{N}\Big)$
$\leq m\Big(r,\frac{F_{N}^{\prime}}{1-F_{N}}\Big)+m\Big(r,\frac{1-F_{N}}{F_{N}}\Big)+N\Big(r,1/F_{N}\Big)$
$\leq m\Big(r,\frac{F_{N}^{\prime}}{1-F_{N}}\Big)+\log 2+m\Big(r,\frac{1}{F_{N}}\Big)+N\Big(r,1/F_{N}\Big)$
$=m\Big(r,\frac{(1-F_{N})^{\prime}}{1-F_{N}}\Big)+T(r,1/F_{N})+\log 2\leq m\Big(r,\frac{(1-F_{N})^{\prime}}{1-F_{N}}\Big)+T(r,F_{N})+\log 2$
$\leq T(r,F_{N})+O\Big(\log^{+}{\frac{N}{1-r}}+\sup_{|z|=(1+r)/2}\log^{+}{\log{|F_{N}|}}\Big),$
@@
At this point the key identity (6.2.5) allows to combine the estimates (6.2.7) and (6.2.9), arriving at the uniform bound
$T(r,p\circ F_{N})=m(r,p\circ F_{N})\leq m\Big(r,\frac{f^{\prime}}{f}\Big)+m\Big(r,\frac{F_{N}}{F_{N}^{\prime}}\Big)$
(6.2.11) $\leq T(r,F_{N})+O\Big(\log^{+}{\frac{N}{1-r}}+\sup_{|z|=(1+r)/2}\log^{+}{\log{|F_{N}|}}\Big).$
@@
Upon doubling the implicit absolute coefficient, plainly for $N\geq 2$ this is equivalent to

### v4

@@ 5.2. Geometry of $\Gamma_{N}$ and a uniform growth estimate of $F_{N}$
The idea of the proof is to use the symmetry of $\mathbf{C}\smallsetminus\mu_{N}$ and the action of the Fuchsian group $\widetilde{\Gamma}_{N}$ to reduce the question to the study of the asymptotic of $\widetilde{F}_{N}$ near the cusp $\tau=i\infty$. We study this asymptotic using the explicit description of the inverse of $\widetilde{F}_{N}$ given in Lemma 5.1.15.
@@
The group $\widetilde{\Gamma}_{N}$ is the free group on $N$ generators given by $\widetilde{t}_{N}$ and its conjugates by powers of $\widetilde{r}_{N}$ in (5.1.3). A fundamental domain $\Omega\subset D(0,1)$ for $\Gamma_{N}$ is the region with $2N$ cusps given by half-integer powers of $\zeta_{N}=\exp(2\pi i/N)$ and bounded by geodesics connecting adjacent cusps.
@@
Since $\widetilde{F}_{N}$ is the universal covering map of $\mathbf{C}\smallsetminus\mu_{N}$, the Fuchsian group $\widetilde{\Gamma}_{N}$ is generated by the stabilizers of the cusps $c$ with $\widetilde{F}_{N}(c)\in\mu_{N}$. If we denote the generator of the stabilizer of $i\infty$ by
@@
$\{z\in D(0,1)\,:\,d(gz,0)\geq d(z,0)\ \text{for all}\ g,g^{-1}\in\{r_{N}^{k}\,t\,r_{N}^{-k}\},k=0,1,\ldots,N-1\}.$
Here, $d$ is the hyperbolic distance in $D(0,1)$. The region in $D(0,1)$ such that $d(gz,0)\geq d(z,0)$ and $d(g^{-1}z,0)\geq d(z,0)$ for $g=r_{N}^{k}\cdot t\cdot r_{N}^{-k}$ is the region bounded by two geodesics starting at $\zeta_{N}^{k}$ going in opposite directions and intersecting the boundary at $\zeta_{N}^{k}e^{\pm i\theta}$ where $c_{N}=2\cot(\theta/2)$. There are exactly $2N$ such arcs corresponding to the $N$ generators and their inverses. In particular, if $\theta<\pi/N$ is too small, the fundamental region will have infinite volume, whereas if $\theta>\pi/N$ is too big, then the Dirichlet domain will only contain at most $N$ cusps. Since $\mathbf{H}/\widetilde{\Gamma}_{N}$ has $N+1$ cusps and $\widetilde{\Gamma}_{N}$ has finite covolume, we must have $c_{N}=2\cot(\pi/2N)$ and these geodesics intersecting at $\zeta_{N}^{k+1/2}$ for $k=0,\ldots,N-1$. ∎
@@
An example of the fundamental domain in Corollary 5.2.4 for $N=3$ is given in Figure 5.2.5, which also includes translates of the domain by elements in $\Omega^{\prime}_{3}$ by words in $\{r_{3},r^{2}_{3},t_{3},t^{-1}_{3}\}$ of length at most $6$. The shading reflects where the absolute value of $|F^{3}_{3}-1|$ is small — it vanishes precisely at the cusps corresponding to $z=1$.
Figure 5.2.5. A (partial) tiling of $D(0,1)$ by a fundamental domain for $\Phi_{3}$, together with a density plot of $|F^{3}_{3}-1|$ which vanishes at $z=1$.
@@
where $\psi$ is the digamma function as in Lemma 5.1.19, and the implicit constant depends on $M_{0}$ but not on $N,M$.
@@
where the implicit constants are uniform in $N$, from which the result follows (using that $|\log{x}|\gg N$). ∎
@@
once $N\gg_{\epsilon,M_{0}}1$, where the implicit constant depends only on $M_{0}$ and $\epsilon$.
@@
If we choose $\varepsilon_{0}:=\epsilon/2$, then for $N\gg_{\epsilon,M_{0}}1$ this lower bound clearly exceeds
$\frac{2N^{2}M(1-\epsilon)}{\pi^{2}}$
as desired. ∎
@@
Set $S(M,N):=\{z\in D(0,1)\,:\,|F_{N}^{N}(z)-1|<e^{-MN}\}$ and $M:=\frac{N}{1-r}$. Since $M\geq 2$, we may take $M_{0}=2$ and $\epsilon=1/2$ in Lemma 5.2.12 and conclude that for $N\gg 1$ (with absolute implicit constant here), we have
@@
By Lemma 5.2.6, the set $\{z\in D(0,1)\,:\,|F_{N}(z)|>e^{M}+1\}$ is contained in $S(M,N)$, which is contained in $D(0,1)\smallsetminus\overline{D(0,r)}$ by the above argument for $N\gg 1$. Thus we conclude that
@@
A more refined bound is proved in Kraus–Roth [KR16, Theorems 1.2 and 1.10]. On the other hand, one can push our method further and prove, with rather more work but uniformly in $N\in\mathbf{N}_{>0}$ and $M\in[1,\infty)$, that the supremum region $|F_{N}|<e^{M}$ is simply connected of conformal radius $1-O(M^{-2}N^{-3})$ from the origin; this is a sharp estimate. But taking for $\varphi$ in Corollary 2.0.5 the pullback of $F_{N}$ by the Riemann map of some such region $|F_{N}|<e^{M}$, and ignoring thus the fine savings from the integrated bound (2.0.7) as opposed to the supremum, would only lead to an $O(N^{4})$ holonomy rank bound in place of our requisite logarithmically inflated bound $O(N^{3}\log{N})$. In the next section we will see how to make the full use of the integrated holonomy bound, and use Nevanlinna’s value distribution theory to supply our final piece of the proof of the unbounded denominators conjecture.
@@ 6. Nevanlinna theory and uniform mean growth near the boundary
for any universal covering map $F:D(0,1)\to\mathbf{C}\smallsetminus\{a_{1},\ldots,a_{N}\}$ based at $F(0)=0$ (see also the discussion in Nevanlinna [Nev70, page 272]), however this is only asymptotically in $r\to 1^{-}$ for given punctures $\{a_{i}\}$ whereas we need a uniformity in both $r$ and $N$. It is at the point (6.2.4) exploiting the small [^3] Precisely, the relevant point of the specific puncture set $\{a_{1},\ldots,a_{N}\}\cup\{\infty\}=\mu_{N}\cup\{\infty\}$ in $\mathbf{P}^{1}\smallsetminus D(0,1)$ is that the degree-$N$ polynomial $\prod_{i=1}^{N}(x-a_{i})\in\mathbf{C}[x]$ has $N^{O(1)}$ coefficients. coefficients of $\prod_{\zeta\in\mu_{N}}(x-\zeta)=x^{N}-1$ that our argument below makes a critical use of the special feature of the target set $\mu_{N}\cup\{\infty\}$ of omitted values.
@@
we have uniformly in $N\in\mathbf{N}_{>0}$ and $r\in(0,1)$ the mean growth bound
@@
with some (effectively computable) absolute implicit constant.
@@ 6.1.1. The Nevanlinna characteristic
where, in general for a meromorphic mapping $f:D(0,1)\to\mathbf{P}^{1}$, we denote by $\mathrm{ord}_{\rho}^{-}(f):=\mathrm{ord}^{+}(1/f)=\max(0,\mathrm{ord}(1/f))$ the pole order (if $\rho$ is a pole, and 0 if $f$ is holomorphic at $\rho$).
@@
For every meromorphic function $f:D(0,1)\to\mathbf{P}^{1}$ regular at 0 (that is: with $f(0)\neq\infty$), and for every $r\in(0,1)$, we have
@@ 6.1.6. The lemma on the logarithmic derivative
The lemma on the logarithmic derivative—a strong explicit form of which is cited in (6.1.14) below—is the centerpiece of Rolf Nevanlinna’s original analytic proof of his second main theorem of value distribution theory. The logarithmic error feature of this sharp upper bound on the proximity function of a logarithmic derivative enables us to derive Theorem 6.0.1 from the relatively crude supremum growth bound in Lemma 5.2.18.
@@
Differentiation in the integrand of (6.1.10) gives a reproducing kernel for our logarithmic derivative as well:
(6.1.11) $\frac{g^{\prime}(z)}{g(z)}=\int_{|w|=R}\frac{2w}{(w-z)^{2}}\log{|g(w)|}\,\mu_{\mathrm{Haar}}(w),\qquad z\in D(0,R).$
@@
$\begin{split}\int_{|z|=r}\Big|\frac{g^{\prime}(z)}{g(z)}\Big|\,\mu_{\mathrm{Haar}}\leq 2R\int_{|z|=r}\int_{|w|=R}|w-z|^{-2}\,\big|\log{|g(w)|}\big|\,\mu_{\mathrm{Haar}}(w)\,\mu_{\mathrm{Haar}}(z)\\ =2R\int_{|w|=R}\Big(\int_{|z|=r}|w-z|^{-2}\,\mu_{\mathrm{Haar}}(z)\Big)\,\big|\log{|g(w)|}\big|\,\mu_{\mathrm{Haar}}(w)\\ =\frac{2R}{R^{2}-r^{2}}\int_{|w|=R}\big|\log{|g(w)|}\big|\,\mu_{\mathrm{Haar}}(w)\\ =\frac{2R}{R^{2}-r^{2}}\Big(m(R,g)+m(R,1/g)\Big)=\frac{4R\,m(R,g)}{R^{2}-r^{2}},\end{split}$
@@
$\begin{split}\int_{|z|=r}\log^{+}{\Big|\frac{g^{\prime}}{g}\Big|}\,\mu_{\mathrm{Haar}}\leq\mu_{\mathrm{Haar}}(E)\log^{+}\Big(\frac{1}{\mu_{\mathrm{Haar}}(E)}\int_{E}\Big|\frac{g^{\prime}(z)}{g(z)}\Big|\,\mu_{\mathrm{Haar}}(z)\Big)\\ \leq\mu_{\mathrm{Haar}}(E)\log^{+}\Big(\frac{1}{\mu_{\mathrm{Haar}}(E)}\int_{|z|=r}\Big|\frac{g^{\prime}(z)}{g(z)}\Big|\,\mu_{\mathrm{Haar}}(z)\Big)\\ \leq\mu_{\mathrm{Haar}}(E)\log^{+}\Big(\int_{|z|=r}\Big|\frac{g^{\prime}(z)}{g(z)}\Big|\,\mu_{\mathrm{Haar}}(z)\Big)+\mu_{\mathrm{Haar}}(E)\log(1/\mu_{\mathrm{Haar}}(E))\\ \leq\log^{+}\int_{|z|=r}\Big|\frac{g^{\prime}(z)}{g(z)}\Big|\,\mu_{\mathrm{Haar}}(z)+\sup_{t\in(0,1]}\big\{t\log{(1/t)}\big\}\\ \leq\log^{+}{\Big\{\frac{4R\,m(R,g)}{R^{2}-r^{2}}\Big\}}+\frac{1}{e}\leq\log^{+}{\Big\{\frac{m(R,g)}{r}\frac{R}{R-r}\Big\}}+\log{2}+\frac{1}{e},\end{split}$
@@
and proved that it is essentially best-possible in form apart for the value of the free numerical constant $5.8501$ (that has since been somewhat further reduced in the literature, see Benbourenane–Korhonen [BK01]). The paper of Hinkkanen [Hin92] and the books of Cherry–Ye [CY01] and Ru [Ru21] discuss the implications to the structure of the error term in Nevanlinna second main theorem, mirroring Osgood and Vojta’s dictionary to Diophantine approximation and comparing to Lang’s conjecture modeled on Khinchin’s theorem.
@@ 6.2.2. Equivalence of bounds for different $p(x)$
(6.2.3) $\begin{split}T(r,p\circ F_{N})=m\Big(r,1+\frac{1}{F_{N}^{N}-1}\Big)\geq m\Big(r,\frac{1}{F_{N}^{N}-1}\Big)-\log{2}\\ =T\Big(r,\frac{1}{F_{N}^{N}-1}\Big)-\log{2}=T(r,F_{N}^{N}-1)-\log{2}\\ \geq T(r,F_{N}^{N})-2\log{2}=N\,T(r,F_{N})-\log{4},\end{split}$
@@ 6.2.8. Two corollaries of the lemma on the logarithmic derivative
$\begin{split}m\Big(r,\frac{f^{\prime}}{f}\Big)\ll\log^{+}{T\Big(\frac{1+r}{2},f\Big)}+\log{\frac{e}{1-r}}\\ =\log^{+}{m\Big(\frac{1+r}{2},1-F_{N}^{N}\Big)}+\log{\frac{e}{1-r}}\\ \ll\log^{+}{m\Big(\frac{1+r}{2},F_{N}^{N}\Big)}+\log{\frac{e}{1-r}}\\ \ll\log^{+}{m\Big(\frac{1+r}{2},F_{N}\Big)}+\log{\frac{N}{1-r}}\\ \ll\sup_{|z|=(1+r)/2}\log^{+}{\log{|F_{N}|}}+\log{\frac{N}{1-r}},\end{split}$
@@
(6.2.13) $\begin{split}m\Big(r,\frac{F_{N}}{F_{N}^{\prime}}\Big)=m\Big(r,\frac{F_{N}^{\prime}}{F_{N}}\Big)+N\Big(r,\frac{F_{N}^{\prime}}{F_{N}}\Big)-N\Big(r,\frac{F_{N}}{F_{N}^{\prime}}\Big)-\log{c(F_{N}^{\prime}/F_{N},0)}\\ =m\Big(r,\frac{F_{N}^{\prime}}{F_{N}}\Big)+N\Big(r,1/F_{N}\Big)-N\Big(r,F_{N}\Big)-N\Big(r,1/F_{N}^{\prime}\Big)+N\Big(r,F_{N}^{\prime}\Big)\\ =m\Big(r,\frac{F_{N}^{\prime}}{F_{N}}\Big)+N\Big(r,1/F_{N}\Big)-N\Big(r,1/F_{N}^{\prime}\Big)=m\Big(r,\frac{F_{N}^{\prime}}{F_{N}}\Big)+N\Big(r,1/F_{N}\Big).\end{split}$
@@
$\begin{split}m\Big(r,\frac{F_{N}}{F_{N}^{\prime}}\Big)=m\Big(r,\frac{F_{N}^{\prime}}{F_{N}}\Big)+N\Big(r,1/F_{N}\Big)\\ \leq m\Big(r,\frac{F_{N}^{\prime}}{1-F_{N}}\Big)+m\Big(r,\frac{1-F_{N}}{F_{N}}\Big)+N\Big(r,1/F_{N}\Big)\\ \leq m\Big(r,\frac{F_{N}^{\prime}}{1-F_{N}}\Big)+\log 2+m\Big(r,\frac{1}{F_{N}}\Big)+N\Big(r,1/F_{N}\Big)\\ =m\Big(r,\frac{(1-F_{N})^{\prime}}{1-F_{N}}\Big)+T(r,1/F_{N})+\log 2\leq m\Big(r,\frac{(1-F_{N})^{\prime}}{1-F_{N}}\Big)+T(r,F_{N})+\log 2\\ \leq T(r,F_{N})+O\Big(\log^{+}{\frac{N}{1-r}}+\sup_{|z|=(1+r)/2}\log^{+}{\log{|F_{N}|}}\Big),\end{split}$
@@
At this point the key identity (6.2.7) allows us to combine the estimates (6.2.10) and (6.2.12), arriving at the uniform bound
(6.2.15) $\begin{split}T(r,p\circ F_{N})=m(r,p\circ F_{N})\leq m\Big(r,\frac{f^{\prime}}{f}\Big)+m\Big(r,\frac{F_{N}}{F_{N}^{\prime}}\Big)\\ \leq T(r,F_{N})+O\Big(\log^{+}{\frac{N}{1-r}}+\sup_{|z|=(1+r)/2}\log^{+}{\log{|F_{N}|}}\Big).\end{split}$
@@
Upon doubling the absolute implicit constant, plainly for $N\geq 2$ this is equivalent to